(1 point) A poll is taken in which 399 out of 575 randomly selected voters indicated their preference for a certain candidate (a) Find a 90% confidence interval for p SPS (b) Find the margin of error for this 90% confidence interval for p (c) without doing any calculations, indicate whether the margin of error is larger or smaller or the same for an 80% confidence interval OA. larger B. smaller C. same
Answer #1

Solution :

Given that,

n = 575

x = 399

Point estimate = sample proportion = \hat p = x / n = 399 / 575 = 0.694

1 - \hat p = 1 - 0.594 = 0.406

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

(b)

Margin of error = E = Z\alpha / 2 * \sqrt((\hat p * (1 - \hat p)) / n)

= 1.645 * (\sqrt((0.694 * 0.406) / 575)

= 0.032

(b)

A 90% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.694 - 0.032 < p < 0.694 + 0.032

0.662 < p < 0.726

(c)

smaller